OpenStudy (anonymous):

Easy, try it out guys, Given 6 different toys of green colour, 5 different toys of blue colour and 4 different toys of red colour, the combination of toys that can be chosen taking atleast one green and one blue toys are ?

OpenStudy (anonymous):

@Ishaan94

OpenStudy (callisto):

how many are you choosing each time?

OpenStudy (pokemon23):

what type of question does this category follow to probability ?

OpenStudy (anonymous):

@Callisto: Any combination ;-)

OpenStudy (anonymous):

@pokemon23: Combinatorics

OpenStudy (pokemon23):

Interesting when do you learn combinatorics?

OpenStudy (callisto):

should i list out all possibility ?

OpenStudy (anonymous):

Is it (11)9! ? heh :$

OpenStudy (callisto):

i think ' the combination of .... are' is not the same as ' how many combinations ..'? :S

OpenStudy (anonymous):

\( (2^6-1)\times (2^5-1)\times 2^4 \)

OpenStudy (mani_jha):

You can do it in two ways: 1)6C1x5C1(4C0+4C1+4C2+4C3+4C4)+6C2x5C1(4C0+4C1+4C2+4C3+4C4)+6C3x5C1(4C0+4C1+4C2+4C3+4C4)+.......+6C6x5C1(4C0+4C1+4C2+4C3+4C4).. =5C1(6C0+6C1+6C2+6C3+...6C6)(4C0+4C1+4.. 2)The total no of ways of choosing atleast one from 6 green balls is (2^6-1) The total no of ways of choosing atleast one from 5 blue balls is (2^5-1) The total no of ways of choosing any from 4 red balls is (2^4) Both will give the same answer! xD